OpenStudy (anonymous):

NEEP HELP MEDAL 1) Suppose f(t)=tlnt, c=2e^2, find (f^-1)|y=c

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

\[f(c)=f(2e^2)=2e^2\ln (2e^2)=2e^2\left(2(\ln 2+\ln e)\right)=4e^2(\ln 2+1)\] Are you supposed to find \(f^{-1}(y)\) when \(y=c\)?

OpenStudy (anonymous):

your supposed to find the inverse of f at y=c I think

OpenStudy (anonymous):

I think what you wrote is right

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

@hartnn @jim_thompson5910 @ganeshie8 need help

OpenStudy (anonymous):

anyone know how to do this?